OpenStudy (anonymous):

I need help with some questions, please help, I will give out blue ribbons!!!!!!!!!!!!!

OpenStudy (anonymous):

Candice bought a pencil box, shown below, to take with her to school. https://app35.studyisland.com/pics/160743pencilboxc.png What is the volume of the pencil box? Candice bought a pencil box, shown below, to take with her to school. *Picture not drawn to scale What is the volume of the pencil box?

OpenStudy (anonymous):

a. 15 7/60 cu in. b. 101 1/2 cu in. c. 40 3/5 cu in. d. 50 3/4 cu in.

OpenStudy (anonymous):

Please, anyone....... HELP!!!

OpenStudy (anonymous):

Anyone!!!! I will give blue ribbons!!!

OpenStudy (saeedd):

volume = l x w x h

OpenStudy (saeedd):

basically lenght times width times height

OpenStudy (saeedd):

so here it would be 9 2/3 X 4 1/5 X 1 1/4

OpenStudy (saeedd):

Do that and tell me

OpenStudy (dayakar):

let length (L) = 9 2/3 = 29/3 in width (W) =41/5 = 21/5 in height (H) = 1 1/4 = 5/4 in volume of the box (V) = L * W *H = 29/ 3 * 21/5 * 5/4 in simplify it u can get the solution